OpenStudy (anonymous):

Subtract the complex expressions, and simplify. (-1.9 + 8.5i) – (-9.5 – 0.4i)

OpenStudy (anonymous):

(-1.9 + 8.5i) – (-9.5 – 0.4i) =-1.9+8.5i+9.5 +0.4i <--- remove the brackets =(-1.9+9.5) + (8.5+0.4)i <--- grouping terms Last step is left for you